OpenStudy (anonymous):

Give an equestion for the line that contains ( 3,2 ) and (5,-2 )

OpenStudy (mertsj):

Slope = (-2-2)/(3-5)=-4/-2=2 So now we know the slope is 2 plug one of the points and the slope into point-slope form.

OpenStudy (mertsj):

Can you do that?

OpenStudy (anonymous):

I Dont Get It

OpenStudy (anonymous):

find slope m = y2 - y1 / x2 - x1 m = -2 - 2 / 5 - 3 m =-4 / 2 m = -2 find b y = mx + b pick any coordinate 2 = -2 (3) + b 2 = -6 + b 8 = b your equation is y = -2x + 8

OpenStudy (mertsj):

Here is point-slope form: y-y1=m(x-x1)

OpenStudy (mertsj):

Ok. Dada will help you, I see.

OpenStudy (anonymous):

lol i showed all the steps so you can learn from it I believe Mertsj is right you should learn how to do it yourself

OpenStudy (anonymous):

How would you grap that

OpenStudy (amistre64):

most basic graphing is to fill in some xs to find the ys to plot

OpenStudy (amistre64):

the equation is a rule for the graph; all it has to be be is followed

OpenStudy (amistre64):

another simpler way is to plot the 2 points given and draw a line between them

OpenStudy (anonymous):

but how would you do that

OpenStudy (amistre64):

the equation itself just defines the line;

OpenStudy (amistre64):

they gave you 2 points right? you know how to plot a point?

OpenStudy (anonymous):